# Evaluating and Mitigating Linguistic Discrimination in Large Language Models

### v1

@@ 1. Introduction
Large language models (LLMs) (Zhao et al., 2023) have attracted considerable public attention, particularly since the emergence of ChatGPT (OpenAI, [n. d.]), which demonstrates remarkable effectiveness in solving diverse natural language processing tasks, including information extraction (Wei et al., 2023), question answering (Tan et al., 2023), and machine translation (Peng et al., 2023). LLM-based chatbots, such as ChatGPT, have seamlessly integrated into the daily routines of many individuals, serving as personal assistants and search engines. Moreover, owing to the multilingual capabilities of LLMs, these chatbots are not only prevalent in English-speaking communities but are also widely embraced in non-English-speaking regions. Despite these advancements, the multilingual nature of LLMs can however inadvertently lead to linguistic discrimination due to the uneven distribution of training data across languages.
@@ 2.3. Large language models
LLMs are pre-trained language models with tens or even thousands of billions (Zhao et al., 2023), such as LLaMA (Touvron et al., [n. d.]) (a collection of models ranging from 7B to 65B parameters) and GPT-3 (Brown et al., 2020)(with 175B parameters). Compared to smaller language models such as GPT-2 (Radford et al., 2019), LLMs not only achieve substantial performance improvements on traditions tasks (Chang et al., 2023), but also exhibit emergent abilities previously unseen in smaller models. There are three typical kinds of emergent abilities exhibited by LLMs, i.e., in-context learning, instruction following, and step-by-step reasoning (Kojima et al., 2022). In-context learning refers to that LLMs can complete a new task without additional training when provided with several samples in the prompt (Dong et al., 2022). Instruction following, achieved through instruction tuning with various tasks, allows LLMs to follow instructions in the prompt without explicit examples (Zeng et al., 2023). Step-by-step reasoning involves LLMs solving complex tasks incrementally using the chain-of-thought (CoT) prompting strategy (Kojima et al., 2022; Wei et al., 2022)
@@ 1 $\mathcal{V}\leftarrow$ encode each response $r\in R$;
3 for each answer $v\in\mathcal{V}$ do
4       $s\leftarrow avgCos(\mathcal{V})$ ;
@@ 8 $r\leftarrow$ response responding to $\hat{v}$;
(8) $avgCos(\mathcal{V})=\frac{\sum_{j\neq k}Cos(v_{k},v_{j})}{|\mathcal{V}|-1}$
@@ 4.3. Results
Figure 11 shows the MJR of four LLMs without or with the help of LDFighter. The values responding to “ori” on x-axis are the result of each LLM without LDFighter, the values responding to “eng” are the result of each LLM with LDFighter where only one language, i.e., English, is used. When the value of $k$ is set to be 3 or above, the MJR of each LLM drops significantly compared to the original MJR. Particularly, the MJR of GPT-3.5, Gemini-pro and Llama2-13b falls straight to 0.0 when using the top three languages, and then remain unchanged as $k$ increases. For Gemma-7b, the MJR first decreases to the lowest point at $k=3$, then rises slightly and stabilizes around 0.1 after $k=18$. When using only English in LDFighter, only Llama2-13b achieves an MJR of 0.0, while the other three LLMs do not reach their optimal MJR. Based on these experimental results, we can answer RQ4 as follows.
Answer to RQ4: LDFighter demonstrates significant effectiveness in reducing the jailbreak rate of harmful questions. Moreover, the use of three languages with the highest CI leads to the lowest MJR for three out of the four LLMs.

### v2

@@ 1. Introduction
Large language models (LLMs) (Zhao et al., 2023b) have attracted considerable public attention, particularly since the emergence of ChatGPT (OpenAI, [n. d.]), which demonstrates remarkable effectiveness in solving diverse natural language processing tasks, including information extraction (Wei et al., 2023), question answering (Tan et al., 2023), and machine translation (Peng et al., 2023). LLM-based chatbots, such as ChatGPT, have seamlessly integrated into the daily routines of many individuals, serving as personal assistants and search engines. Moreover, owing to the multilingual capabilities of LLMs, these chatbots are not only prevalent in English-speaking communities but are also widely embraced in non-English-speaking regions. Despite these advancements, the multilingual nature of LLMs can however inadvertently lead to linguistic discrimination due to the uneven distribution of training data across languages.
@@ 2.3. Large language models
LLMs are pre-trained language models with tens or even thousands of billions (Zhao et al., 2023b), such as LLaMA (Touvron et al., [n. d.]) (a collection of models ranging from 7B to 65B parameters) and GPT-3 (Brown et al., 2020)(with 175B parameters). Compared to smaller language models such as GPT-2 (Radford et al., 2019), LLMs not only achieve substantial performance improvements on traditions tasks (Chang et al., 2023), but also exhibit emergent abilities previously unseen in smaller models. There are three typical kinds of emergent abilities exhibited by LLMs, i.e., in-context learning, instruction following, and step-by-step reasoning (Kojima et al., 2022). In-context learning refers to that LLMs can complete a new task without additional training when provided with several samples in the prompt (Dong et al., 2022). Instruction following, achieved through instruction tuning with various tasks, allows LLMs to follow instructions in the prompt without explicit examples (Zeng et al., 2023). Step-by-step reasoning involves LLMs solving complex tasks incrementally using the chain-of-thought (CoT) prompting strategy (Kojima et al., 2022; Wei et al., 2022)
@@ 1 $\mathcal{V}\leftarrow$ encode each response $r\in R$;
3 for each answer $v_{k}\in\mathcal{V}$ do
4       $s\leftarrow avgCos(\mathcal{V},v_{k})$ ;
@@ 8 $r\leftarrow$ response responding to $\hat{v}$;
(8) $avgCos(\mathcal{V},v_{k})=\frac{\sum_{j\neq k}Cos(v_{k},v_{j})}{|\mathcal{V}|-1}$
@@ 4.3. Results
Figure 11 shows the MJR of four LLMs without or with the help of LDFighter. The values responding to “ori” on x-axis are the result of each LLM without LDFighter, the values responding to “eng” are the result of each LLM with LDFighter where only one language, i.e., English, is used. When the value of $k$ is set to be 3 or above, the MJR of each LLM drops significantly compared to the original MJR. Particularly, the MJR of GPT-3.5, Gemini-pro and Llama2-13b falls straight to 0.0 when using the top three languages, and then remain unchanged as $k$ increases. For Gemma-7b, the MJR first decreases to the lowest point at $k=3$, then rises slightly and stabilizes around 0.1 after $k=18$. When using only English in LDFighter, only Llama2-13b achieves an MJR of 0.0, while the other three LLMs do not reach their optimal MJR.
Table 1. ASR of GCG adversarial samples on LDFighter-equipped Llama2-7b
Top-k 3 4 5 6 7 8 9 10
ASR 13% 8% 8% 6% 5% 4% 2% 2%
In addition to addressing inherent multilingual challenges, we further conduct a simple experiment to assess LDFighter’s effectiveness in defending against state-of-the-art jailbreaking attacks. Specifically, we randomly selected adversarial prompts targeting Llama2-7b from the dataset created by CASPER (Zhao et al., 2023a), generated using the GCG approach (Zou et al., [n. d.]). Subsequently, LDFighter is applied to Llama2-7b to counter these prompts. To determine the optimal language selection strategy, we consider the top-k languages based on their average CI score as revealed in RQ3. Varying k from 3 to 10, we record the corresponding attack success rates (ASR), as detailed in Table 1. Note that all selected adversarial prompts are in English and validated, resulting in an initial ASR of 100% on Llama2-7b. Notably, the integration of LDFighter leads to a significant reduction in ASR. Even with just the top 3 languages (including English), the ASR plummets to 13% from the baseline 100%. Furthermore, utilizing the top 9 languages, the ASR drops to a mere 2%. These results underscore the remarkable potential of LDFighter in thwarting state-of-the-art jailbreaking attacks. Based on these experimental results, we can answer RQ4 as follows.
Answer to RQ4: LDFighter demonstrates significant effectiveness in reducing the jailbreak rate of both harmful questions and adversarial prompts generated by GCG. Moreover, the use of three languages with the highest CI leads to the lowest MJR for three out of the four LLMs.